# Formatting Instructions For NeurIPS 2024

### v1

@@ 5.1.1 Semantic Latent for Downstream Tasks
The representation ${\boldsymbol{\mathbf{z}}}_{\mathrm{sem}}$ should be general and transferable  [14] for downstream tasks. We measure the quality of learned representations in classification tasks. To ensure their universality, we assess the quality of the learned representations on various datasets by different classifiers. Specifically, Following [48], we train a classifier on labelled test sets for each ParamReL model. We allocate $80\%$ of the dataset for training a classifier and reserve the remaining $20\%$ for testing purposes. The performance on the test set is evaluated based on AUROC. This process is conducted in a 5-fold cross-validation manner, with the results reported as mean metrics ± one standard deviation. The results are shown in Table 2. Higher AUROC would suggest that the learned features ${\boldsymbol{\mathbf{z}}}_{\mathrm{sem}}$ contain more information about data ${\boldsymbol{\mathbf{x}}}$. In addition to assessing the representation quality, we also compare the image reconstruction ability against baselines. From Table 2, we can conclude that VAE-based models still produce blurry reconstructions, while diffusion-based and parameter-based models can build near-exact reconstructions. Refer to Figure 9 in Appendix G for reconstructed binary images.
@@ 5.2.3 Smooth Representation Learning for Latent Interpolation
Latent space interpolation  [15, 17] can validate the smoothness, continuity, and semantic coherence in the learned representation learning space of generative models. Generally, we can embed two samples into the corresponding latent space, and interpolating between latent variables sampled from the corresponding space yields interpolated representations. Their reconstruction learned by generative models reveals the semantic richness of the learned space. The specific latent interpolation process is detailed in Appendix E.
@@ 5.2.4 Disentanglement
To facilitate a thorough and unbiased quantitative assessment of disentanglement, we employ two metrics: a prediction-based indicator: Disentanglement, Completeness, and Informativeness (DCI) [11], and an intervention-based criterion: Total AUROC Difference (TAD) [52]. Meanwhile, we report the reconstruction quality in Appendix G and conclude that ParamReL can achieve near-exact reconstruction on Celeba (Figure 12), 3DShapes (Figure 13) and CIFAR10 (Figure 14).

### v3

@@ 5.1.1 Semantic Latent for Downstream Tasks
The representation ${\boldsymbol{\mathbf{z}}}_{\mathrm{sem}}$ should be general and transferable  franceschi2019unsupervised for downstream tasks. We measure the quality of learned representations in classification tasks. To ensure their universality, we assess the quality of the learned representations on various datasets by different classifiers. Specifically, Following xiao2023trading , we train a classifier on labelled test sets for each ParamReL model. We allocate $80\%$ of the dataset for training a classifier and reserve the remaining $20\%$ for testing purposes. The performance on the test set is evaluated based on AUROC. This process is conducted in a 5-fold cross-validation manner, with the results reported as mean metrics ± one standard deviation. The results are shown in Table 2. Higher AUROC would suggest that the learned features ${\boldsymbol{\mathbf{z}}}_{\mathrm{sem}}$ contain more information about data ${\boldsymbol{\mathbf{x}}}$. In addition to assessing the representation quality, we also compare the image reconstruction ability against baselines. From Table 2, we can conclude that VAE-based models still produce blurry reconstructions, while diffusion-based and parameter-based models can build near-exact reconstructions. Refer to Figure 9 in Appendix G for reconstructed binary images.
@@ 5.2.3 Smooth Representation Learning for Latent Interpolation
Latent space interpolation  goodfellow2014generative ; higgins2017beta can validate the smoothness, continuity, and semantic coherence in the learned representation learning space of generative models. Generally, we can embed two samples into the corresponding latent space, and interpolating between latent variables sampled from the corresponding space yields interpolated representations. Their reconstruction learned by generative models reveals the semantic richness of the learned space. The specific latent interpolation process is detailed in Appendix E.
@@ 5.2.4 Disentanglement
To facilitate a thorough and unbiased quantitative assessment of disentanglement, we employ two metrics: a prediction-based indicator: Disentanglement, Completeness, and Informativeness (DCI) eastwood2018framework , and an intervention-based criterion: Total AUROC Difference (TAD) yeats2022nashae . Meanwhile, we report the reconstruction quality in Appendix G and conclude that ParamReL can achieve near-exact reconstruction on Celeba (Figure 12), 3DShapes (Figure 13) and CIFAR10 (Figure 14).